# Time to Retrain? Detecting Concept Drifts in Machine Learning Systems

### v1

@@ IV-D Clustering and Sampling for Label Requisition
However, when there are no high-density cores due to randomly generated features (SEA and Sine datasets in RQ1 and RQ2), DBSCAN can form too many or only a single cluster. We address this via K-means clustering [77, 75]. The initial centroids are randomized with a count of $\frac{w}{100}$ to ensure at most 1% of the dataset requires labels. The K-means algorithm computes the Euclidean distance and assigns points closest to the centroids to their respective clusters where the mean of each cluster becomes the new centroids. This process is repeated until convergence yielding the clusters surrounding the centroids.
@@
When inputs are images or text (as in RQ3), we also leverage K-means for clustering after pre-processing the input. For images, the features are the pixels of the image with some caveats. First, the pixels are in grayscale and normalized to have a value between 0 and 1. Second, the size of each image needs to be very small (no more than 28x28 pixels). Datasets without these characteristics may need feature reduction and grayscale conversion. We can then compute distance treating each pixel as a feature. When the inputs are texts, we follow the original pre-processing steps to clean, trim, vectorize, and tokenize the input texts [78]. This results in the features being fixed-sized sparse feature vectors with numbered values representing the encoded vocabulary that can be used to compute distances representing similarities between texts.
Stratified random sampling data points. Via DBSCAN or K-means, the clusters from $W$ can be treated as strata to be sampled for labeling.
Expert labeling. Experts will label a subset of this sample similar to how SOTA requires experts to label the selected data points.
Rationale behind clustering. SOTA semi-supervised methods request labels based on metrics of the target data/model. For example, ECHO selects data points for labeling based on the classifier model’s confidence. Depending on the metric, SOTA may request labels for 0-100% testing data which causes the detection process to fail or be very inefficient. On the other hand, simple random label requests are inefficient and costly since we risk labeling unnecessary points. Hence, clustering is a compromise that directs the manual labeling process to the most representative points in each sample via stratified sampling.
Discussion on how CDSeer addresses industry needs. The clustering process is meant to address several industry needs. First, data points are selected so experts can label multiple points in one session. Second, the data points summarize the current data, minimizing the effort. Third, not all selected data points need to be labeled, giving the expert some choice. This allows for the flexibility needed in the industry compared to SOTA.
@@ IV-E Propagation and Retraining the Inspector
All manually labeled data points are managed by memory $M$. $M$ is a fixed-sized queue of the newest data points with true labels available where the oldest points are automatically removed when the pre-defined size of $M$ is exceeded. Since the true labels (in $M$) are scarce, we increase their number via semi-supervised learning. To do this, we combine $M$ and $W$ and apply semi-supervised learning so true labels spread from $M$ to $W$ to get $(W,TrueLabels)$ where all points have labels. We select $LabelSpreading$ in sklearn [79], a graph inference-based algorithm that is robust to noises and can accurately spread very few labeled points to a large number of unlabeled points. This method constructs a complete weighted graph between all data points (labeled and unlabeled) where the weights are the distance between the points. Labels would spread from labeled to unlabeled data points iteratively until convergence where a higher weight slows spreading. If an unlabeled data point is assigned multiple different labels, the most frequently assigned label is selected as the final label for the point. Then, we retrain $InspectorModel$ using $(W,TrueLabels)$.
@@ V Evaluation Setup
This section describes the evaluation setup including computing resources, evaluation metrics, datasets, and candidate models included in this paper.
V-A Computing Environment
In RQ1 and RQ2, we experimented on a Windows computer with a Core i7-8650U processor and 32 GiB of RAM. In RQ3, to speed up the CDD experiments, the evaluations are performed on a system running Ubuntu 20.04 equipped with an AMD Ryzen 16-core processor, 256 GiB of RAM, and an NVIDIA GeForce RTX 3090 GPU.
V-B Evaluation Datasets
@@
We select eight datasets for our evaluation as shown in Table I. These publicly available datasets are collected from previous works on concept drift detection [26, 6] which consisted of three synthetic datasets (Sine, SEA, FM) and four real-world datasets (ELEC, NOAA, LMR, GCT). Also, we include one internal, real-world run-time data collected from the deployment of congestion prediction at a telecommunication company, i.e., ROUTER. In our use case, the datasets can be divided into two groups: with GT and without GT concept drift. The following datasets contain known concept drifts: Sine, SEA, FM, LMR, and ROUTER. Of these datasets, Sine, SEA, and ROUTER 100% do not have data drift. Details of the dataset are as follows:
@@
Electricity (ELEC) [26]: This dataset contains data points for two months between 1996 and 1998. Each data point shows demands in different locations over a period of 30 minutes. The label identifies the change in the price related to a moving average of the last 24 hours.
@@
NOAA [81] is a subset of the original NOAA dataset that contains 50 years of weather data from Offutt Air Force Base in Bellevue, Nebraska. The label represents whether rain was observed on each day.
@@
Google cluster trace (GCT) [6]: first released by Google in 2011 [11]. This dataset contains the cleaned and under-sampled version of the original dataset detailing the trace data of a production cluster. The features are of the first five minutes of monitored data of the job, and the labels reflect whether the job eventually failed. This dataset is extremely imbalanced, hence, we down-sampled the non-failure class at a 3:1 ratio, leading to less than 3.20% data points that belong to the failure class.
@@
Fashion-MNIST (FM) [82] contains images showing individual articles of clothing at low resolution (28 by 28 pixels). We reduce the number of classes to two. If the class is 0, 2, 3, 4, 6 it is a Top and the label is 0. Otherwise, it is 1. We create a synthetic concept drift at 4,000 as follows. From index 4,000, if the point’s original classes are 4 or 6 (now both are 0), it is switched to 1. In this case, it roughly translated to 366 items that have their labels changed. The concept drift detection is performed on the 10,000 images in the testing dataset.
@@
Large Movie Review (LMR) [83] contains class balance movie reviews from the Internet Movie Database. We introduce concept drift at index 2,000 where the classes are reversed.
V-C Subject ML Models Trained Using the Datasets
@@
We train three types of online models for the evaluation. In RQ1 and RQ2, the online models are random forest classifiers (RFC). For RQ3, two models are trained: a three-layer NN for the FM dataset and a five-layer NN for the LMR dataset.
@@
We train the online models in RQ1 and RQ2 using the RFC implementation in scikit-learn [84]. The training data consists of the first 500/1,000 data points of the corresponding datasets including labels. The remainder of the datasets are converted into a time series without labels for testing. Hence, all experiments are conducted online where concept drift detection is active for the entire time series.
In RQ3, we evaluate the generality of CDSeer, i.e., to what extent CDSeer can detect concept drifts in different ML models as a model-agnostic approach. We trained two NNs implemented by Keras Sequential models [85, 86] using the code and data from TensorFlow [87, 78].
The first NN contains three layers for classifications of clothing images in the Fashion-MNIST dataset. The first layer flattens the input. Then, the second layer (128 nodes) learns the clothing features. Lastly, the third layer (10 nodes) outputs the classification results. This neural network is trained on the first 60,000 data points of this FM dataset.
@@
For the textual LMR dataset, the neural network for sentiment classification consists of five layers. The first layer takes the integer-encoded text and looks up an embedding vector for each word. The third layer converts the sequence to a fixed length. The last layer is densely connected with a single output node. The second and fourth rounds are drop-out layers to reduce overfitting. This neural network is trained on the first 25,000 data points of this LMR dataset.
@@ V-D Evaluation Metrics
  • False positive: The number of concept drifts detected by one CDD that are incorrect and do not match with the ground-truth concept drifts.
@@
VI Evaluation Results
@@
Motivation. Motivated by our industry needs, we want to examine the performance of the SOTA concept-drift detection techniques on both open-source and proprietary datasets. We want to learn what their limitations are, what we can learn from them, and whether we can use the insights to devise better ideas for concept drift detection. First, we want to learn how they perform on the core functionality such as concept drift detection and the resulting accuracy. Then we want to learn how they perform on practical concerns of the concept drift detectors such as the false positive rate and the amount of labels needed.
Method. We answer this RQ using one supervised and one semi-supervised CDD approach. We set up a well-known SOTA semi-supervised concept drift detection method for comparison, i.e., ECHO [8]. It contains performance improvements over its predecessor, SAND [39], and is effective in detecting concept drift. We acquired their official Java implementation [88]. We run ECHO using their default configuration only changing the warm-up window size (to -S 1000). The default parameters are ideally their best configuration, and this change has the effect of allowing ECHO to cover up to at least 70% of the data points of each dataset. This choice comes after experimenting with various configurations, which end up with ECHO crashing when less than 30% of datasets are executed, trivializing the comparison effort. We confirmed that the results are at least on par with that observed in the original papers of ECHO and SAND [8, 39] or prior papers using them as a baseline [9, 12] on equivalent datasets.
For supervised CDD approaches, we include the Page-Hinkley test (PHT) [27] implementation in scikit-multiflow [75]. PHT is a well-known supervised CDD approach commonly used in previous work [18, 42, 57]. PHT is a variant of the Cumulative Sum [27], another well-known CDD used as the basis for concept drift detection in ECHO. As we use PHT as a supervised CDD, we allow its access to 100% of the ground-truth labels in the test datasets.
These methods are evaluated on the datasets described in Section V-B. As described in Section V-C, to evaluate PHT, we train a random forest classifier on the first 1,000 data points of the dataset with true labels to create the base model. The rest of the dataset is the test dataset. Once concept drift is detected, the base model is retrained using the most recent 1,000 data points with true labels.
TABLE II: The performance of three concept-drift detection (CDD) approaches. PHT is a supervised CDD, which requires 100% manual labeling and is unrealistic in practice. ECHO and Our Prophet are semi-supervised CDD and are more practical.
@@
Google Cloud 96.91 97.37 69.2 53.8 100.0 96.74 1.3 75.0 91.7    97.33 17.6 35.7 0.8
Results. Table II shows the results of applying PHT (a supervised CDD) and ECHO (the SOTA semi-supervised) to detect concept drifts in the six datasets. We also show the model accuracy of the original online model on the datasets for comparison. It is expected that with the help of CDD, the model accuracy will increase. Also, it is expected that the supervised CDD PHT would have a better performance than the semi-supervised CDD ECHO.
Our evaluation results show that PHT and ECHO can improve the overall model accuracy, i.e., PHT contributes to a higher model accuracy for five of the six datasets and ECHO leads to a higher or comparable model accuracy for two of the six datasets. The limited improvements of PHT and ECHO are expected for two reasons. First, the datasets contain a relatively small number of concept drifts that span a small percentage of data points. Second, these datasets contain class imbalance, hence, wrong classifications for minority classes are not as impactful. In such cases, adapting to the new concepts could cause slight decreases in accuracy. In reality, online models that take streaming data accumulated in weeks and months will have a much larger benefit from concept drift detection.
In terms of concept drift detection, our evaluation results confirm that the supervised CDD PHT achieves 100% precision in the three datasets with ground-truth concept drifts, i.e., ROUTER, SEA, and Sine. PHT fails to detect some concept drifts in ROUTER and Sine, i.e., the recall is 50%. A lower recall is expected considering the conservative design of PHT, which favors precision over recall. As a semi-supervised CDD, ECHO’s precision is lower and more unstable (ranging from 1.3% to 100%) on the six datasets compared to PHT. The recall of ECHO is comparable to that of PHT, ranging from 50% to 100%.
Another important metric of CDD is the amount of ground-truth labels required, i.e., expert labeling needs to be performed manually. A highly practical CDD is expected to have a low percentage of manual labels required. As PHT is supervised, it requires 100% of the data to be manually labeled. The semi-supervised CDD ECHO requires significantly fewer labels. However, our evaluation results find that the percentage of labels required by ECHO is unstable with a large range across different datasets, as low as 0.03% for Sine, yet an excessive amount of labeling is required for the other datasets (48.1% to 100%).
@@
Motivation. As the SOTA CDD techniques do not satisfy our industry needs, we propose CDSeer including two key innovations, i.e., clustering data points and spreading true labels to reduce the excessive labeling required while improving model accuracy. In this RQ, we examine the performance of CDSeer. First, we want to evaluate the performance of CDSeer in concept drift detection using the default configuration. In addition, as CDSeer is configurable, to investigate the impacts of CDSeer’s configurations, we examine the performance of CDSeer under different configurations.
@@
- Using CDSeer’s default configuration. We set up CDSeer to run on datasets with known concept drifts, i.e., ROUTER, SEA, and Sine. For each of the three datasets, we use the first 1,000 data points to train a random forest classifier to be the online model. In this part, we set the windows and memory sizes to our default configuration, The default window size is 1,000 and the default memory size is 15.
- Using different configurations of CDSeer. We set off to investigate how the different settings of window and memory size would impact the performance of CDSeer. In CDSeer, the window and memory size used to train the inspector model have a significant impact on the performance of concept drift detection. The window influences the clustering process and influences which data points are labeled and stored in the memory. After the labels are spread from memory, the labeled window is also the training data for the inspector model. To evaluate the impact, we vary the window and the memory size used to train the inspector model and observe the resulting performance. We experiment with memory with queue sizes 10 and 15. For the window size, we try two values of 500 and 1000, where the latter is the same size as the training data used to train the online models. Each pair of memory and window size will affect the retraining process of the inspector model, changing its prediction and the resulting concept drift detections.
As CDSeer integrates a supervised CDD to detect changes in the distribution of the prediction error between the original and the inspector models (Algorithm 1), we use PHT as the integrated supervised CDD in this experiment. Similar to RQ1, we evaluate model accuracy, precision, recall, and percentage of labels required.

### v2

@@ IV-D1 Clustering Data Stream
However, when there are no high-density cores due to randomly generated features, DBSCAN can form too many or only a single cluster. We address this via K-means clustering [77, 75]. The initial centroids are randomized with a count of $\frac{w}{100}$ to ensure at most 1% of the dataset requires labels. The K-means algorithm computes the Euclidean distance and assigns points closest to the centroids to their respective clusters where the mean of each cluster becomes the new centroids. This process is repeated until convergence yielding the clusters surrounding the centroids.
@@
When inputs are images or text, we also leverage K-means for clustering after pre-processing the input. For images, the features are the pixels of the image with some caveats. First, the pixels are in grayscale and normalized to have a value between 0 and 1. Second, the size of each image needs to be very small (no more than 28x28 pixels). Datasets without these characteristics may need feature reduction and grayscale conversion. We can then compute distance treating each pixel as a feature. When the inputs are texts, we follow the original pre-processing steps to clean, trim, vectorize, and tokenize the input texts [78]. This results in the features being fixed-sized sparse feature vectors with numbered values representing the encoded vocabulary that can be used to compute distances representing similarities between texts.
IV-D2 Sampling from Clusters
From the previous step, $W$ is clustered into several $clusters$ of points. Then, one point from each cluster is sampled for labeling. The points within the same cluster are meant to be interchangeable.
IV-D3 Labeling by Experts
Then, experts will label the sample similar to how they label requests in SOTA methods, with two improvements: (1) the expert can label only a subset of the sample, and (2) the expert can exchange the sampled point with another from the same cluster if desired.
Rationale behind clustering. We arrive at clustering from our observations of how the current SOTA works. SOTA semi-supervised methods request labels based on metrics of the target data/model. For example, ECHO selects data points for labeling based on the classifier model’s confidence. Depending on the metric, SOTA may request labels for 0-100% of the testing data, which can cause the detection process to fail or be highly inefficient. On the other hand, simple random label requests are inefficient and costly, as we risk labeling unnecessary points. Hence, clustering is a compromise that directs the manual labeling process to the most representative points in each window via stratified sampling.
Discussion on how CDSeer addresses industry needs. Clustering is meant to address several industry needs. First, data points are selected so experts can label multiple points in a single session. Second, the data points summarize the current data, preventing efforts spent labeling highly similar points. Third, not all selected data points need to be labeled, and data points within the same strata are interchangeable, allowing experts to label the points that require the least effort. This approach offers greater flexibility needed in the industry compared to SOTA.
@@ IV-E Propagation and Retraining the Inspector
[…]
@@ V Evaluation Setup
We conduct an evaluation to answer three research questions (RQs). This section describes the evaluation metrics, datasets, and candidate models included in the evalution.
RQ3-Generality: Can CDSeer detect concept drift in pre-trained models?
V-A Evaluation Datasets
@@
We select eight datasets for our evaluation (Table I). Seven of which are publicly available collected from previous works on concept drift detection [26, 6] which consisted of three synthetic datasets (Sine, SEA, FM) and four real-world datasets (ELEC, NOAA, LMR, GCT). Also, we include one internal, real-world run-time dataset collected from the deployment of congestion prediction at a telecommunication company, i.e., ROUTER. In our use case, the datasets can be divided into two groups: with ground-truth (GT) and without GT concept drift. The following datasets contain known concept drifts: Sine, SEA, FM, LMR, and ROUTER. Of these datasets, Sine, SEA, and ROUTER 100% do not have data drift. Details of the dataset are as follows:
@@
Electricity (ELEC) [26]: This dataset includes data points representing 30-minute demand data from various locations, covering two months between 1996 and 1998. The label indicates the price change relative to the moving average of the past 24 hours. Data points representing price increases comprised 42.5% of the dataset.
@@
NOAA [81]: A subset of the original NOAA dataset containing 50 years of weather data from Offutt Air Force Base in Bellevue, Nebraska. The label represents whether rain was observed on each day, where rainy days comprised 31.4% of the data.
@@
Google cluster trace (GCT) [6]: This dataset contains the cleaned and under-sampled version of the original production cluster trace data released by Google in 2011 [11]. The features represent the first five minutes of monitored job data, while the labels indicate whether the job ultimately failed. This dataset is extremely imbalanced. The non-failure class consisted of less than 3.2% of the dataset even after a 3:1 down-sampling.
@@
Fashion-MNIST (FM) [82]: A dataset contains images showing individual articles of clothing at low resolution (28 by 28 pixels). We reduce the number of classes to two. If the class is 0, 2, 3, 4, 6 it is a Top and the label is 0. Otherwise, it is 1. After reduction, this is a balanced dataset. We create a synthetic concept drift at 4,000 as follows. From index 4,000, if the point’s original classes are 4 or 6 (now both are 0), it is switched to 1. In this case, it roughly translated to 366 items with their labels changed. The concept drift detection is performed on the 10,000 images in the testing dataset.
@@
Large Movie Review (LMR) [83]: A balanced movie reviews dataset from the Internet Movie Database. We introduce concept drift at index 2,000 where the classes are reversed.
V-B Subject ML Models Trained Using the Datasets
@@
In RQ1 and RQ2, the online models are random forest classifiers. For RQ3, two models are trained: a three-layer NN for the FM dataset and a five-layer NN for the LMR dataset.
@@
We train the online models in RQ1 and RQ2 using the random forest classifiers implementation in scikit-learn [84]. The training data consists of the first 500/1,000 data points of the corresponding datasets including labels. The remainder of the datasets are converted into a time series without labels for testing. Hence, all experiments are conducted online where concept drift detection is active for the entire time series.
@@
In RQ3, we evaluate the generality of CDSeer, i.e., to what extent CDSeer can detect concept drifts in different ML models as a model-agnostic approach. We trained two NNs implemented by Keras Sequential models [85, 86] using the code and data from TensorFlow [87, 78]. The first NN contains three layers for classifications of clothing images in the Fashion-MNIST dataset. The first layer flattens the input. Then, the second layer (128 nodes) learns the clothing features. Lastly, the third layer (10 nodes) outputs the classification results. This neural network is trained on the first 60,000 data points of this FM dataset. The second NN consists of five layers for sentimental classification of the textual LMR dataset. The first layer takes the integer-encoded text and looks up an embedding vector for each word. The third layer converts the sequence to a fixed length. The last layer is densely connected with a single output node. The second and fourth rounds are drop-out layers to reduce overfitting. This neural network is trained on the first 25,000 data points of this LMR dataset.
@@ V-C Evaluation Metrics
  • False positive: The number of concept drifts detected by one CDD that are incorrect and do not match with the GT concept drifts.
@@
VI Evaluation
This section provides the details for each RQ, including the motivation, method, and results.
@@ VI Evaluation
Motivation. Motivated by our industry needs, we want to examine the performance of the SOTA concept drift detection techniques on both open-source and proprietary datasets. We want to learn what their limitations are, what we can learn from them, and whether we can use the insights to devise better ideas for concept drift detection. First, we want to learn how they perform on core functionality such as concept drift detection and the resulting accuracy. Then we want to learn how they perform on practical concerns of the CDDs such as the false positive rate and the amount of labels needed.
Method. We answer this RQ by evaluating a supervised CDD (i.e., PHT [27]) and a semi-supervised CDD (i.e., ECHO [8]) on the first six datasets in Table I. ECHO is a well-known SOTA semi-supervised CDD that improves upon its predecessor [39]. We run its official Java implementation [88] using their default configuration, only changing the warm-up window size (to -S 1000). The default parameters are ideally their best configuration, and this change enables ECHO to cover at least 70% of the data points of each dataset. This choice comes after experimenting with various configurations, which end up with ECHO crashing when less than 30% of datasets are executed, trivializing the comparison effort. We provide the true labels for all of ECHO’s label requests. We confirmed that the results are at least on par with that observed in ECHO’s papers [8, 39] or works using them as baselines [9, 12] on equivalent datasets. PHT is a supervised CDD commonly used in previous work [18, 42, 57]. PHT is a variant of the Cumulative Sum [27], another well-known CDD used as the basis for concept drift detection in ECHO. For our experiment, we use the PHT implementation in scikit-multiflow [75]. As we use PHT as a supervised CDD, we allow its access to 100% of the GT labels in the test datasets.
These methods are evaluated on the datasets described in Section V-A. As described in Section V-B, to evaluate PHT, we train a random forest classifier on the first 1,000 data points of the dataset with true labels to create the base model. The rest of the dataset is the test dataset. Once concept drift is detected, the base model is retrained using the most recent 1,000 data points with true labels.
TABLE II: The performance of three CDDs: PHT (supervised), ECHO (semi-supervised), and CDSeer (semi-supervised).
@@
GCT 96.91 97.37 69.2 53.8 100.0 96.74 1.3 75.0 91.7    97.33 17.6 35.7 0.8
Results. Table II shows the results of applying PHT and ECHO for concept drift detection across six datasets. We also show the model accuracy of the original online model on the datasets for comparison. It is expected that with the help of a CDD, the model accuracy will increase. Also, it is expected that the supervised CDD PHT outperforms the semi-supervised CDD ECHO.
Our evaluation results show that PHT and ECHO can improve the overall model accuracy, i.e., PHT contributes to a higher model accuracy for five of the six datasets and ECHO leads to a higher or comparable model accuracy for two of the six datasets. The limited improvements of PHT and ECHO are expected for two reasons. First, the datasets contain a relatively small number of concept drifts that span a small percentage of data points. Second, these datasets contain class imbalance, hence, wrong classifications for minority classes are not as impactful. In such cases, adapting to the new concepts could cause slight decreases in accuracy. In practice, online models that process streaming data accumulated over weeks or months would greatly benefit from concept drift detection.
In terms of concept drift detection, our evaluation results confirm that the supervised CDD PHT achieves 100% precision in the three datasets with GT concept drifts, i.e., ROUTER, SEA, and Sine. PHT fails to detect some concept drifts in ROUTER and Sine, i.e., the recall is 50%. A lower recall is expected considering the conservative design of PHT, which favors precision over recall. As a semi-supervised CDD, ECHO’s precision is lower and more unstable (ranging from 1.3% to 100%) on the six datasets compared to PHT. The recall of ECHO is comparable to that of PHT, ranging from 50% to 100%.
Another important metric of CDDs is the amount of GT labels required, i.e., expert labeling needs to be performed manually. A highly practical CDD is expected to have a low percentage of manual labels required. As PHT is supervised, it requires 100% of the data to be manually labeled. The semi-supervised CDD ECHO requires significantly fewer labels. However, our evaluation results find that the percentage of labels required by ECHO is unstable with a large range across different datasets, as low as 0.03% for Sine, yet an excessive amount of labeling is required for the other datasets (48.1% to 100%).
@@
Current SOTA either demands impractically large amounts of labeled data or is incompatible with pre-trained industry models.
@@
Motivation. Since the SOTA CDDs do not satisfy industry needs, we propose CDSeer including two key innovations, i.e., clustering data points and spreading true labels to reduce the excessive labeling required while improving model accuracy. In this RQ, we want to examine the performance of CDSeer in two steps. First, we evaluate the performance of CDSeer in concept drift detection using the default configuration. Second, we want to examine the performance of CDSeer under different window and memory sizes. These two parameters are configurable and significantly influence CDSeer’s effectiveness in concept drift detection. The window determines which data points are labeled and stored in the memory, while the memory determines the training data labels for the inspector model.
@@
- Using CDSeer’s default configuration. We run CDSeer on the first six datasets in Table I. For each dataset, we use the first 1,000 data points to train a random forest classifier to be the online model. In this part, we set the windows and memory sizes to our default configuration. The default window size is 1,000 and the default memory size is 15.
- Using different configurations of CDSeer. We run CDSeer on the three datasets with known concept drifts, i.e., ROUTER, SEA, and Sine. We vary the window and the memory size used to train the inspector model and observe the resulting impact on performance. We experiment with memory with queue sizes 10 and 15. For the window size, we try two values of 500 and 1,000, where the latter is the same size as the training data used to train the online models. Each pair of memory and window size will affect the retraining process of the inspector model, changing its prediction and the resulting concept drift detected.
For both steps, we use PHT as the integrated supervised CDD to detect changes in the prediction error distribution between the original and the inspector models (Algorithm 1). Similar to RQ1, we provide the CDD with all labels requested and evaluate model accuracy, precision, recall, and percentage of labels required.